# Beyond Discretization: Learning the Optimal Solution Path

### v2

@@ 1 Introduction
$\boldsymbol{\theta}^{*}(\lambda)\in\arg\min_{\boldsymbol{\theta}\in\mathbb{R}^{d}}h(\boldsymbol{\theta},\lambda),\quad\lambda\in\Lambda,$ (1)
where $\Lambda$ is an arbitrary set of parameters indexing the problems. (We assume Equation 1 admits an optimal solution for each $\lambda\in\Lambda$.) In such applications, we often seek to compute the entire solution path $\{\boldsymbol{\theta}^{*}(\lambda):\lambda\in\Lambda\}$ in order to present experts with a portfolio of possible solutions to compare and assess tradeoffs.
As an example, consider the $p$-norm fair facility location problem (Gupta et al., 2023), which minimizes facility opening costs while incorporating $l_{p}$-regularization to promote fairness across socioeconomic groups. Since there is no obvious choice for $p$, we might prefer computing solutions for all $p$ and allowing experts to (qualitatively) assess the resulting solutions. Many other applications entail navigating similar tradeoffs, including upweighting the minority class in binary classification to tradeoff Type I and II errors or aggregating features to increase interpretability at the expense of accuracy. In each case, we seek the entire solution path because selecting the “best” solution requires weighing a variety of criteria, some of which may be qualitative. These settings differ from classical hyperparameter tuning where there is a clear auxiliary criterion (like out-of-sample performance), and it would be enough to identify the single $\lambda^{*}\in\Lambda$ such that $\boldsymbol{\theta}^{*}(\lambda^{*})$ optimizes this criteria.
The most common approach to learning the entire solution path is discretization: discretize $\Lambda$, solve Equation 1 at each grid point, and interpolate the resulting solutions. With enough grid points, interpolated solutions are approximately optimal along the entire path. Several authors have sought to determine the minimal the number of discretization points needed to achieve a target level of accuracy, usually under minimal assumptions on $h(\boldsymbol{\theta},\lambda)$. Giesen et al. (2010) considers convex optimization problems over the unit simplex when $\Lambda\subseteq\mathbb{R}$ and show that learning the solution path to accuracy $\epsilon$ requires at least $O(1/\epsilon)$ grid points. Giesen et al. (2012) consider the case where $h(\boldsymbol{\theta},\lambda)$ is concave in $\lambda$ and $\Lambda\subseteq\mathbb{R}$ and show only $O(1/\sqrt{\epsilon})$ points are needed. More recently, Ndiaye et al. (2019) relate the required number of points to the regularity of $h(\boldsymbol{\theta},\lambda)$, arguing that if $h$ is uniformly convex of order $d$, one requires $O(1/\sqrt[d]{\epsilon})$ points.
@@ 1.1 Our Contributions
We propose a novel, simple algorithmic procedure to learn the solution path that applies to an arbitrary set $\Lambda$. The key idea is to replace the family of problems in (1) with a single stochastic optimization problem. This stochastic optimization problem depends on two user-specified components: a distribution $\mathbb{P}_{\lambda}$ over values of $\lambda\in\Lambda$ and a collection of basis functions $\boldsymbol{\Phi}_{j}(\cdot):\Lambda\to\mathbb{R}^{d}$, $j=1,\ldots,p$. We then seek to approximate $\boldsymbol{\theta}^{*}(\lambda)$ as a linear combination of basis functions, $\boldsymbol{\Phi}_{1:p}(\lambda)\hat{\boldsymbol{\beta}}$, where $\boldsymbol{\Phi}_{1:p}:=\left[\boldsymbol{\Phi}_{1}\ \boldsymbol{\Phi}_{2}\ \cdots\ \boldsymbol{\Phi}_{p}\right]$ and $\hat{\boldsymbol{\beta}}$ is an approximate solution to
$\min_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\ \mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\boldsymbol{\Phi}_{1:p}(\tilde{\lambda})\boldsymbol{\beta},\tilde{\lambda})\right].$ (2)
@@
Figure 1: Learning Solution Path of Weighted Binary Classification. See Section 6 for setup. We compare our method with $p=5,7,9$ Legendre polynomials as our basis with uniform discretization. Orange line is our adaptive method (c.f. Algorithm 1). Vertical lines indicate when new basis functions are added.
@@
    When using constant step-size SGD to solve Equation 2, we prove that the uniform error of our learned path $\boldsymbol{\Phi}_{1:p}(\lambda)\hat{\boldsymbol{\beta}}$ to the true path $\boldsymbol{\theta}^{*}(\lambda)$ exhibits linear convergence to an irreducible constant that is proportional to the expressiveness of the basis (Theorem 3.4). This behavior is already visible in Figure 1.
The proof of this result utilizes ideas from the convergence of SGD under various “growth conditions” of the gradient (Bottou et al., 2018; Nguyen et al., 2018; Vaswani et al., 2019; Liu et al., 2024; Bertsekas, 1996). See Khaled and Richtárik (2020) for a summary and comparison of these various conditions. Our contribution to this literature is to relate the expressiveness of the basis $\boldsymbol{\Phi}_{1:p}(\lambda)$ to a relaxed, weak-growth condition of Gower et al. (2019, Lemma 2.4). Indeed, we show that under some assumptions, Equation 2 always satisfies this relaxed weak-growth condition, and if the solution path lies in the span of the basis, it satisfies the weak-growth condition of Vaswani et al. (2019). This allows us to leverage results from those works to establish Theorem 3.4.
In special cases, we can leverage a priori knowledge of the structure of $\boldsymbol{\theta}^{*}(\lambda)$ to prove stronger results. For example, suppose $\Lambda=[-1,1]$, and we use Legendre polynomials as our basis. We prove that
@@
    If the solution path $\boldsymbol{\theta}^{*}(\lambda)$ is $\nu$-times differentiable, then using $p=O\left(\epsilon^{\frac{1}{2(1-\nu)}}\right)$ polynomials ensures that after $T=O\left(\epsilon^{\frac{1}{1-\nu}}\log(1/\epsilon)\right)$ gradient calls, we obtain an $\epsilon$-approximation to the solution path (Theorem 4.3).
For comparison, Ndiaye et al. (2019) implies that when $h(\boldsymbol{\theta},\lambda)$ is strongly-convex, discretization requires at least $O(\epsilon^{-1/2})$ points. Hence, even if the optimization problem at each grid point can be solved with $1$ gradient evaluation, our approach requires asymptotically fewer evaluations whenever $\nu>3$. If $\nu$ is large, the savings is substantive.
@@
    If the solution path $\boldsymbol{\theta}^{*}(\lambda)$ is analytic ($\nu=\infty$), then using $p=O(\log(1/\epsilon))$ basis polynomials ensures that after $T=O(\log^{2}(1/\epsilon)\log\log(1/\epsilon))$ iterations, we obtain an $\epsilon$-approximation to the solution path (Theorem 4.5).
This is almost exponentially less work than discretization.
@@
1:  Initialize: $\boldsymbol{\Phi}_{1:0}(\cdot)\leftarrow[\ ]$ and $\hat{\boldsymbol{\beta}}_{0}\leftarrow[\ ]$
2:  Return: A sequence of coefficients: $\hat{\boldsymbol{\beta}}_{1},\hat{\boldsymbol{\beta}}_{2},\ldots$
@@
$\boldsymbol{\Phi}_{1:p}(\cdot)\leftarrow\left[\boldsymbol{\Phi}_{1:{p-1}}(\cdot)\quad\boldsymbol{\Phi}_{p}(\cdot)\right]$
@@
6:   $[\hat{\boldsymbol{\beta}}_{p-1},0]\in\mathbb{R}^{p}$, to obtain
$\hat{\boldsymbol{\beta}}_{p}~{}\approx~{}\arg\min_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\ \mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\boldsymbol{\Phi}_{1:p}(\tilde{\lambda})\boldsymbol{\beta},\tilde{\lambda})\right].$
@@ 1.2 Other Related Work
There are some works that consider the total computational work to learn a solution path, typically focusing on a specific choice of $h(\cdot,\cdot)$. See, e.g., Tibshirani (1996); Cortes and Vapnik (1995); Rosset (2004); Rosset and Zhu (2007); Friedman et al. (2010); Osborne et al. (2000); Hastie et al. (2004); Efron et al. (2004); Mairal and Yu (2012); Tibshirani and Taylor (2011); Ndiaye and Takeuchi (2021) for results on LASSO and regularized SVM; Park and Hastie (2007); Bao et al. (2019) for results on $L1$-regularization. Liu and Grigas (2023) study a regularization path setting from an ordinary differential equations perspective using second-order methods. Overall, our work is similar to this larger stream in that we focus on total computational cost, but, unlike this stream, we develop a general-purpose framework with minimal assumptions on $h(\cdot,\cdot)$.
@@ 2 Model Setup and Preliminaries
There are constants $0<\mu\leq L$ such that, for all $\lambda\in\Lambda$, $h(\cdot,\lambda)$ is $\mu$-strongly convex and $L$-smooth, i.e., for all $\boldsymbol{\theta},\bar{\boldsymbol{\theta}}\in\mathbb{R}^{d}$,
$\frac{\mu}{2}\|\boldsymbol{\theta}-\bar{\boldsymbol{\theta}}\|^{2}~{}\leq~{}h(\boldsymbol{\theta})-h(\bar{\boldsymbol{\theta}})-\nabla h(\bar{\boldsymbol{\theta}})^{\top}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})~{}\leq~{}\frac{L}{2}\|\boldsymbol{\theta}-\bar{\boldsymbol{\theta}}\|^{2}.$
For any candidate solution path $\boldsymbol{\theta}(\cdot)$, we define the solution path error of $\boldsymbol{\theta}(\cdot)$ by
$\epsilon_{\rm sp}(\boldsymbol{\theta}(\cdot)):=\sup_{\lambda\in\Lambda}\left\{h(\boldsymbol{\theta}(\lambda),\lambda)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda)\right\}.$
An $\epsilon$-solution path is a solution path $\boldsymbol{\theta}(\cdot)$ such that $\epsilon_{\rm sp}(\boldsymbol{\theta}(\cdot))<\epsilon$. Finally, given any vector of coefficients $\boldsymbol{\beta}$ and basis functions $\boldsymbol{\Phi}_{1:p}(\cdot)$, we define $\epsilon_{\rm sp}(\boldsymbol{\beta}):=\epsilon_{\rm sp}(\boldsymbol{\Phi}_{1:p}(\cdot)\boldsymbol{\beta})$ to be the solution path error of $\boldsymbol{\Phi}_{1:p}(\cdot)\boldsymbol{\beta}$.
As mentioned, our method depends on two user-chosen parameters: a distribution $\mathbb{P}_{\lambda}$ over $\lambda\in\Lambda$ and a series of basis functions $\boldsymbol{\Phi}_{j}:\Lambda\mapsto\mathbb{R}^{d}$, for $j=1,\ldots,p$. We require some minor assumptions on these choices:
@@
    It is easy to compute $\boldsymbol{\Phi}_{j}(\lambda)\in\mathbb{R}^{d}$ for any $\lambda\in\Lambda$, and $1\leq j\leq p$.
@@
    There does not exist $\boldsymbol{\beta}\in\mathbb{R}^{p}$ with $\boldsymbol{\beta}\neq 0$ such that $\boldsymbol{\Phi}_{1:p}(\tilde{\lambda})\boldsymbol{\beta}=\boldsymbol{0}$ holds $\mathbb{P}_{\lambda}$-almost everywhere.
@@
Given $\boldsymbol{\Phi}_{1:p}$, we define the minimal solution path error for this basis by
$\epsilon_{\rm sp}^{\ast}:=\inf_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\epsilon_{\rm sp}(\boldsymbol{\beta}).$
@@
$C$ $~{}:=~{}\sup_{\lambda\in\Lambda}\sigma_{\max}\left(\boldsymbol{\Phi}(\lambda)^{\top}\boldsymbol{\Phi}(\lambda)\right),$ (3)
$c$ $~{}:=~{}\sigma_{\min}\left\{\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\boldsymbol{\Phi}(\tilde{\lambda})^{\top}\boldsymbol{\Phi}(\tilde{\lambda})\right]\right\}.$
In words, $C$ is a uniform bound on the largest eigenvalue of the positive semidefinite matrix $\boldsymbol{\Phi}(\lambda)^{\top}\boldsymbol{\Phi}(\lambda)$, and $c$ is the smallest eigenvalue of the corresponding expected matrix. By construction, $0\leq c\leq C$. Under Assumption 2.2, both constants are strictly positive.
@@
We can now state our first key result which relates the suboptimality of a feasible solution $\boldsymbol{\beta}$ in Equation 2 to its solution path error. Define
$\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}~{}\in~{}\arg\min_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\ \mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta},\tilde{\lambda})\right].$
@@
Under Assumptions 2.2 and 2.1, for any $\boldsymbol{\beta}\in\mathbb{R}^{p}$, we have
$\epsilon_{\rm sp}(\boldsymbol{\beta})~{}\leq~{}2CL\|\boldsymbol{\beta}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}+\left(\frac{8CL}{c\mu}\right)\epsilon_{\rm sp}^{\ast}.$
This bound decomposes into two terms, one proportional to $\|\boldsymbol{\beta}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}$, which represents the suboptimality of $\boldsymbol{\beta}$ and the other proportional to $\epsilon_{\rm sp}^{\ast}$, which measures the maximal expressiveness of the basis. By solving Equation 2 to greater accuracy, we can drive down the first term, but we will not affect the second term. To reduce the second term, we must add basis functions to obtain a better quality approximation.

### v3

@@ 1 INTRODUCTION
$\bm{\theta}^{*}(\lambda)\in\arg\min_{\bm{\theta}\in\mathbb{R}^{d}}h(\bm{\theta},\lambda),\quad\lambda\in\Lambda,$ (1)
where $\Lambda$ is an arbitrary set of parameters indexing the problems. (We assume Equation 1 admits an optimal solution for each $\lambda\in\Lambda$.) In such applications, we often seek to compute the entire solution path $\{\bm{\theta}^{*}(\lambda):\lambda\in\Lambda\}$ in order to present experts with a portfolio of possible solutions to compare and assess tradeoffs.
As an example, consider the $p$-norm fair facility location problem (Gupta et al., 2023), which minimizes facility opening costs while incorporating $l_{p}$-regularization to promote fairness across socioeconomic groups. Since there is no obvious choice for $p$, we might prefer computing solutions for all $p$ and allowing experts to (qualitatively) assess the resulting solutions. Many other applications entail navigating similar tradeoffs, including upweighting the minority class in binary classification to tradeoff Type I and II errors or aggregating features to increase interpretability at the expense of accuracy. In each case, we seek the entire solution path because selecting the “best” solution requires weighing a variety of criteria, some of which may be qualitative. These settings differ from classical hyperparameter tuning where there is a clear auxiliary criterion (like out-of-sample performance), and it would be enough to identify the single $\lambda^{*}\in\Lambda$ such that $\bm{\theta}^{*}(\lambda^{*})$ optimizes this criteria.
The most common approach to learning the entire solution path is discretization: discretize $\Lambda$, solve Equation 1 at each grid point, and interpolate the resulting solutions. With enough grid points, interpolated solutions are approximately optimal along the entire path. Several authors have sought to determine the minimal the number of discretization points needed to achieve a target level of accuracy, usually under minimal assumptions on $h(\bm{\theta},\lambda)$. Giesen et al. (2012a) considers convex optimization problems over the unit simplex when $\Lambda\subseteq\mathbb{R}$ and show that learning the solution path to accuracy $\epsilon$ requires at least $O(1/\epsilon)$ grid points. Giesen et al. (2012b) consider the case where $h(\bm{\theta},\lambda)$ is concave in $\lambda$ and $\Lambda\subseteq\mathbb{R}$ and show only $O(1/\sqrt{\epsilon})$ points are needed. More recently, Ndiaye et al. (2019) relate the required number of points to the regularity of $h(\bm{\theta},\lambda)$, arguing that if $h$ is uniformly convex of order $d$, one requires $O(1/\sqrt[d]{\epsilon})$ points.
@@ 1.1 Our Contributions
We propose a novel, simple algorithmic procedure to learn the solution path that applies to an arbitrary set $\Lambda$. The key idea is to replace the family of problems in (1) with a single stochastic optimization problem. This stochastic optimization problem depends on two user-specified components: a distribution $\mathbb{P}_{\lambda}$ over values of $\lambda\in\Lambda$ and a collection of basis functions $\bm{\Phi}_{j}(\cdot):\Lambda\to\mathbb{R}^{d}$, $j=1,\ldots,p$. We then seek to approximate $\bm{\theta}^{*}(\lambda)$ as a linear combination of basis functions, $\bm{\Phi}_{1:p}(\lambda)\hat{\bm{\beta}}$, where $\bm{\Phi}_{1:p}:=\left[\bm{\Phi}_{1}\ \bm{\Phi}_{2}\ \cdots\ \bm{\Phi}_{p}\right]$ and $\hat{\bm{\beta}}$ is an approximate solution to
$\min_{\bm{\beta}\in\mathbb{R}^{p}}\ \mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\bm{\Phi}_{1:p}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})\right].$ (2)
@@
Figure 1: Learning Solution Path of Weighted Binary Classification. See Section 6 for setup. We compare our method with $q=5,7,9$ Legendre polynomials as our basis with uniform discretization. Orange line is our adaptive method (c.f. Algorithm 1). Vertical lines indicate when new basis functions are added.
@@
    When using constant step-size SGD to solve Equation 2, we prove that the uniform error of our learned path $\bm{\Phi}_{1:p}(\lambda)\hat{\bm{\beta}}$ to the true path $\bm{\theta}^{*}(\lambda)$ exhibits linear convergence to an irreducible constant that is proportional to the expressiveness of the basis (Theorem 3.4). This behavior is already visible in Figure 1.
The proof of this result utilizes ideas from the convergence of SGD under various “growth conditions” of the gradient (Bottou et al., 2018; Nguyen et al., 2018; Vaswani et al., 2019; Liu et al., 2024; Bertsekas, 1996). See Khaled and Richtárik (2020) for a summary and comparison of these various conditions. Our contribution to this literature is to relate the expressiveness of the basis $\bm{\Phi}_{1:p}(\lambda)$ to a relaxed, weak-growth condition of Gower et al. (2019, Lemma 2.4). Indeed, we show that under some assumptions, Equation 2 always satisfies this relaxed weak-growth condition, and if the solution path lies in the span of the basis, it satisfies the weak-growth condition of Vaswani et al. (2019). This allows us to leverage results from those works to establish Theorem 3.4.
In special cases, we can leverage a priori knowledge of the structure of $\bm{\theta}^{*}(\lambda)$ to prove stronger results. For example, suppose $\Lambda=[-1,1]$, and we use Legendre polynomials as our basis. We prove that
@@
    If the solution path $\bm{\theta}^{*}(\lambda)$ is $\nu$-times differentiable, then using $p=O\left(\epsilon^{\frac{1}{2(1-\nu)}}\right)$ polynomials ensures that after $T=O\left(\epsilon^{\frac{1}{1-\nu}}\log(1/\epsilon)\right)$ gradient calls, we obtain an $\epsilon$-approximation to the solution path (Theorem 4.3).
For comparison, Ndiaye et al. (2019) implies that when $h(\bm{\theta},\lambda)$ is strongly-convex, discretization requires at least $O(\epsilon^{-1/2})$ points. Hence, even if the optimization problem at each grid point can be solved with $1$ gradient evaluation, our approach requires asymptotically fewer evaluations whenever $\nu>3$. If $\nu$ is large, the savings is substantive.
@@
    If the solution path $\bm{\theta}^{*}(\lambda)$ is analytic ($\nu=\infty$), then using $p=O(\log(1/\epsilon))$ basis polynomials ensures that after $T=O(\log^{2}(1/\epsilon)\log\log(1/\epsilon))$ iterations, we obtain an $\epsilon$-approximation to the solution path (Theorem 4.5).
This is almost exponentially less work than discretization for unidimensional hyperparameter space.
@@
1:  Initialize: $\bm{\Phi}_{1:0}(\cdot)\leftarrow[\ ]$ and $\hat{\bm{\beta}}_{0}\leftarrow[\ ]$
2:  Return: A sequence of coefficients: $\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2},\ldots$
@@
$\bm{\Phi}_{1:p}(\cdot)\leftarrow\left[\bm{\Phi}_{1:{p-1}}(\cdot)\quad\bm{\Phi}_{p}(\cdot)\right]$
@@
6:   $[\hat{\bm{\beta}}_{p-1},0]\in\mathbb{R}^{p}$, to obtain
$\hat{\bm{\beta}}_{p}~{}\approx~{}\arg\min_{\bm{\beta}\in\mathbb{R}^{p}}\ \mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\bm{\Phi}_{1:p}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})\right].$
@@ 1.2 Other Related Work
There is a rich literature on specialized methods for computing the solution path under regularization, i.e., where $\lambda\in\mathbb{R}_{+}$ is the weight on a convex regularizer. These methods generally employ a path-following algorithm. See, for example, Rosset (2004); Friedman et al. (2010); Park and Hastie (2007) for cases where $\bm{\theta}^{*}(\lambda)$ is smooth, and Rosset and Zhu (2007) for when it is piecewise linear. The LASSO, or $\ell_{1}$ regularized regression, has received special emphasis (Osborne et al., 2000; Efron et al., 2004; Tibshirani and Taylor, 2011), as has support vector machines (Hastie et al., 2004) and certain structured regularizers (Bao et al., 2019). Liu and Grigas (2023) adopts a slightly different perspective than this existing literature, using ordinary differential equations to motivate a second-order method to compute the path. Overall, like our work, these works generally consider the total computational time to compute the path, not just the number of discretization points.
However, our work differs from this literature in two important ways. First, our proposed method applies to general $\lambda\in\Lambda$ which may be multidimensional, whereas the previous literature largely focuses on the case of $\lambda\in\mathbb{R}_{+}$. This limitation of the previous literature is perhaps fundamental as path-following algorithms do not generalize easily to multidimensional settings. Second, we treat a general function $h(\cdot,\cdot)$, not just the case of regularization. Despite our generality, our paper performs a relatively precise computational analysis by bounding total computation time in terms of the number of gradient calls, i.e., we treat the gradient as a block-box oracle instead of a more complicated black-box consisting of minimizing the objective $h(\cdot,\cdot)$ at fixed parameter values. Furthermore, our numerical experiments go beyond simple regularization examples that are common in the literature by considering the upweighing of a minority class in classification and weighting different components of a multi-objective optimization problem. Thus, our method applies more broadly than the aforementioned works.
@@ 2 MODEL SETUP AND PRELIMINARIES
There are constants $0<\mu\leq L$ such that, for all $\lambda\in\Lambda$, $h(\cdot,\lambda)$ is $\mu$-strongly convex and $L$-smooth, i.e., for all $\bm{\theta},\bar{\bm{\theta}}\in\mathbb{R}^{d}$,
$\frac{\mu}{2}\|\bm{\theta}-\bar{\bm{\theta}}\|^{2}~{}\leq~{}h(\bm{\theta})-h(\bar{\bm{\theta}})-\nabla h(\bar{\bm{\theta}})^{\top}(\bm{\theta}-\bar{\bm{\theta}})~{}\leq~{}\frac{L}{2}\|\bm{\theta}-\bar{\bm{\theta}}\|^{2}.$
For any candidate solution path $\bm{\theta}(\cdot)$, we define the solution path error of $\bm{\theta}(\cdot)$ by
$\epsilon_{\rm sp}(\bm{\theta}(\cdot)):=\sup_{\lambda\in\Lambda}\left\{h(\bm{\theta}(\lambda),\lambda)-h(\bm{\theta}^{*}(\lambda),\lambda)\right\}.$
An $\epsilon$-solution path is a solution path $\bm{\theta}(\cdot)$ such that $\epsilon_{\rm sp}(\bm{\theta}(\cdot))<\epsilon$. Finally, given any vector of coefficients $\bm{\beta}$ and basis functions $\bm{\Phi}_{1:p}(\cdot)$, we define $\epsilon_{\rm sp}(\bm{\beta}):=\epsilon_{\rm sp}(\bm{\Phi}_{1:p}(\cdot)\bm{\beta})$ to be the solution path error of $\bm{\Phi}_{1:p}(\cdot)\bm{\beta}$.
As mentioned, our method depends on two user-chosen parameters: a distribution $\mathbb{P}_{\lambda}$ over $\lambda\in\Lambda$ and a series of basis functions $\bm{\Phi}_{j}:\Lambda\mapsto\mathbb{R}^{d}$, for $j=1,\ldots,p$. We require some minor assumptions on these choices:
@@
    It is easy to compute $\bm{\Phi}_{j}(\lambda)\in\mathbb{R}^{d}$ for any $\lambda\in\Lambda$, and $1\leq j\leq p$.
@@
    There does not exist $\bm{\beta}\in\mathbb{R}^{p}$ with $\bm{\beta}\neq 0$ such that $\bm{\Phi}_{1:p}(\tilde{\lambda})\bm{\beta}=\bm{0}$ holds $\mathbb{P}_{\lambda}$-almost everywhere.
@@
Given $\bm{\Phi}_{1:p}$, we define the minimal solution path error for this basis by
$\epsilon_{\rm sp}^{\ast}:=\inf_{\bm{\beta}\in\mathbb{R}^{p}}\epsilon_{\rm sp}(\bm{\beta}).$
@@
$C$ $~{}:=~{}\sup_{\lambda\in\Lambda}\sigma_{\max}\left(\bm{\Phi}(\lambda)^{\top}\bm{\Phi}(\lambda)\right),$ (3)
$c$ $~{}:=~{}\sigma_{\min}\left\{\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\bm{\Phi}(\tilde{\lambda})^{\top}\bm{\Phi}(\tilde{\lambda})\right]\right\}.$
In words, $C$ is a uniform bound on the largest eigenvalue of the positive semidefinite matrix $\bm{\Phi}(\lambda)^{\top}\bm{\Phi}(\lambda)$, and $c$ is the smallest eigenvalue of the corresponding expected matrix. By construction, $0\leq c\leq C$. Under Assumption 2.2, both constants are strictly positive.
@@
We can now state our first key result which relates the suboptimality of a feasible solution $\bm{\beta}$ in Equation 2 to its solution path error. Define
$\bm{\beta}_{\mathrm{avg}}^{\ast}~{}\in~{}\arg\min_{\bm{\beta}\in\mathbb{R}^{p}}\ \mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})\right].$
@@
Under Assumptions 2.2 and 2.1, for any $\bm{\beta}\in\mathbb{R}^{p}$, we have
$\epsilon_{\rm sp}(\bm{\beta})~{}\leq~{}2CL\|\bm{\beta}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}+\left(\frac{8CL}{c\mu}\right)\epsilon_{\rm sp}^{\ast}.$
This bound decomposes into two terms, one proportional to $\|\bm{\beta}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}$, which represents the suboptimality of $\bm{\beta}$ and the other proportional to $\epsilon_{\rm sp}^{\ast}$, which measures the maximal expressiveness of the basis. By solving Equation 2 to greater accuracy, we can drive down the first term, but we will not affect the second term. To reduce the second term, we must add basis functions to obtain a better quality approximation.